OpenStudy (anonymous):

Find the exact solutions and simplify as much as possible for 5x^2+8=12x

OpenStudy (cwrw238):

5x^2 - 12x + 8 = 0 this will not factor so use quadratic formula x = [-b +/- sqrt((b^2 - 4ac) ] / 2a here a=5,b=-12 and c = 8

OpenStudy (anonymous):

So, how does i come in to play?

OpenStudy (anonymous):

So for instance, when I use the calculator I get 6/5+(2/5)i, 6/5-(2/5)i

OpenStudy (cwrw238):

x = [-(-12) + sqrt(144-4*5*8) ] / 10 = 12 + sqrt(-16) ------------ 10 = 12/10 + (4/10) i = 6/5 + (2/5)i - other root is calculated in similar way

OpenStudy (anonymous):

But what is the i?

OpenStudy (cwrw238):

i = sqrt(-1) - the roots are complex numbers a + bi is complex number where a = real part and bi is imaginary part there is no real square root of a negative number

OpenStudy (anonymous):

ahhhhhh ok

OpenStudy (anonymous):

Thank you very much!

OpenStudy (cwrw238):

yw